### v1

# On Generating Monolithic and Model Reconciling Explanations in Probabilistic Scenarios

Abstract
Explanation generation frameworks aim to make AI systems’ decisions transparent and understandable to human users. However, generating explanations in uncertain environments characterized by incomplete information and probabilistic models remains a significant challenge. In this paper, we propose a novel framework for generating probabilistic monolithic explanations and model reconciling explanations. Monolithic explanations provide self-contained reasons for an explanandum without considering the agent receiving the explanation, while model reconciling explanations account for the knowledge of the agent receiving the explanation. For monolithic explanations, our approach integrates uncertainty by utilizing probabilistic logic to increase the probability of the explanandum. For model reconciling explanations, we propose a framework that extends the logic-based variant of the model reconciliation problem to account for probabilistic human models, where the goal is to find explanations that increase the probability of the explanandum while minimizing conflicts between the explanation and the probabilistic human model. We introduce explanatory gain and explanatory power as quantitative metrics to assess the quality of these explanations. Further, we present algorithms that exploit the duality between minimal correction sets and minimal unsatisfiable sets to efficiently compute both types of explanations in probabilistic contexts. Extensive experimental evaluations on various benchmarks demonstrate the effectiveness and scalability of our approach in generating explanations under uncertainty.
1 Introduction
The rapid integration of artificial intelligence (AI) into critical and everyday applications has magnified the importance of not just achieving high-performance metrics but also ensuring that AI decisions are transparent, interpretable, and, above all, trustworthy. This imperative has given rise to the field of explainable AI (XAI), which seeks to make AI systems’ workings comprehensible to their human users (?). XAI endeavors to demystify the often opaque processes of AI, providing insights into the reasoning behind decisions and actions. This transparency is not just a matter of ethical AI design but a practical necessity for enhancing user trust, facilitating user decision-making, and ensuring the accountability of AI systems.
In the domain of machine learning (ML), significant strides have been made towards enhancing the explainability of algorithms. Researchers have sought to categorize ML algorithms according to various dimensions of explainability (?), improve the transparency of existing algorithms (?, ?), and even propose new algorithms that balance accuracy with increased explainability (?, ?). These efforts underscore a growing recognition within the ML community of the critical role that explainability plays in the deployment of AI systems (?).
Parallel to advancements in ML, the automated planning community has adopted a focused approach to generating explanations for plans produced by AI (planning) agents, which led to the inception of explainable AI planning (XAIP) (?). Predominantly, XAIP research focuses on the explanation generation problem, which involves identifying explanations for plans that, when conveyed to human users, help them understand and accept the agent’s proposed actions (?, ?). A noteworthy direction within this space is the model reconciliation problem (MRP) (?, ?, ?, ?), aimed at aligning a user’s model with that of an AI agent through the provision of explanations, especially when discrepancies in their understanding of a planning problem lead to confusion or misinterpretation of the agent’s decisions. We will refer to these explanations as model reconciling explanations. In contrast, we refer to explanations that do not account for the user’s model as monolithic explanations.
Although MRP tackles essential facets of explainability, it often operates under the presumption that the AI agent has a deterministic grasp of the human model – a scenario that may not always align with the complexities of real-world interactions characterized by uncertainty about human knowledge. Indeed, this gap highlights a general challenge in explanation generation from AI agents: Their operation within realms of incomplete information and probabilistic decision-making models. Traditional explanation methods, which rely on deterministic knowledge, falter under these conditions, unable to represent the uncertain nature of the agent’s knowledge adequately.
To bridge this gap in explainability, the first part of the paper introduces an approach that integrates uncertainty into the explanation generation framework. In particular, using (propositional) probabilistic logic as our underlying mechanism for modeling uncertainty, we propose a framework, where given a belief base $\mathcal{B}$, which is a weighted knowledge base, and an explanandum $\varphi$, the goal is to find a probabilistic monolithic explanation that increases the probability of the explanandum, that is, the probability that the explanandum is true. We introduce the concepts of explanatory gain and explanatory power, which serve as metrics for evaluating the effectiveness of probabilistic monolithic explanations. Furthermore, we extend our framework from logic-based model reconciliation problems (L-MRPs) (?, ?, ?), focusing on scenarios beset with uncertainty in the agent’s perception of the human model. Specifically, given a knowledge base $\mathrm{KB}_{\alpha}$ of an agent, an explanandum $\varphi$ entailed by $\mathrm{KB}_{\alpha}$, and a human belief base $\mathcal{B}_{h}$, the goal is to find a probabilistic model reconciling explanation that not only increases the explanandum’s probability for $\mathcal{B}_{h}$ but also decreases the probability of conflicts between the explanation and $\mathcal{B}_{h}$.
In the second part of the paper, we describe algorithms for computing both probabilistic monolithic and model reconciling explanations, leveraging the conceptual duality between minimal correction sets (MCSes) and minimal unsatisfiable sets (MUSes). These algorithms, adapted from our prior work in ? (?), integrate a weighted maximum satisfiability procedure for computing probabilities. Our experimental evaluation across various benchmarks underscores the practicality and applicability of our algorithms, demonstrating their capability to generate probabilistic explanations within a propositional logic framework.
In summary, our contributions are the following:
  • $\bullet$
    We propose a novel framework for generating probabilistic monolithic explanations given a belief base and an explanandum. Central to our framework are the concepts of explanatory gain and explanatory power, metrics designed to quantitatively assess the quality and effectiveness of probabilistic explanations.
  • $\bullet$
    We present a probabilistic framework for L-MRP to tackle scenarios characterized by uncertainty in the agent’s understanding of the human model.
  • $\bullet$
    We describe algorithms for computing probabilistic monolithic and model reconciling explanations, using the duality of MCSes and MUSes together with a weighted maximum satisfiability process.
  • $\bullet$
    Through a series of benchmark evaluations, we demonstrate the efficacy of our proposed algorithms in generating probabilistic explanations.
The paper is structured as follows. In Section 2, we provide the necessary background knowledge. In Section 3, we describe our explanation generation framework for monolithic and model reconciliation probabilistic explanations. In Section 4, we present algorithms for computing classical and probabilistic explanations, and experimentally evaluate them on a set of benchmarks in Section 5. Finally, in Section 6, we discuss related work, and conclude the paper in Section 7.
2 Background
In this section, we provide some background for propositional logic, the hitting set duality between minimal unsatisfiable and minimal correction sets, modeling uncertainty in propositional logic, and the model reconciliation problem.
2.1 Propositional Logic
Let $\mathcal{L}$ be a propositional language built from a finite set of atomic variables $\mathcal{V}=\{a,b,c,\ldots\}$. A possible world is a truth-value assignment to each variable $\omega:\mathcal{V}\mapsto\{T,F\}$, where $T$ and $F$ denote truth and falsity respectively. The set of all possible worlds of $\mathcal{L}$ is denoted by $\Omega$. The simplest formulae in $\mathcal{L}$ are atoms: Individual variables that may be true or false in a given possible world. More complex formulae are recursively constructed from atoms using the classical logical connectives. A model of a formula is a possible world in which the formula is satisfied (i.e., evaluates to true). A knowledge base $\mathrm{KB}$ is a set of formulae. If there exists at least one possible world $\omega$ that satisfies all formulae in $\mathrm{KB}$, then $\mathrm{KB}$ is consistent, otherwise we say that $\mathrm{KB}$ is inconsistent. We use $\models$ to denote the classical entailment relation and say that a (consistent) $\mathrm{KB}$ entails a formula $\varphi$, expressed as $\mathrm{KB}\models\varphi$, if and only if every model of $\mathrm{KB}$ is also a model of $\varphi$, or equivalently, if $\mathrm{KB}\cup\{\lnot\varphi\}$ is inconsistent. Unless stated otherwise, it is assumed that all formulae are expressed in conjunctive normal form (CNF). [^1] A CNF formula is a conjunction of clauses, where each clause is a disjunction of literals. A literal is either an atom or its negation. This is not a restrictive requirement, since any propositional formula can be transformed into a CNF representation.
Given a knowledge base $\mathrm{KB}$ and a formula $\varphi$, called the explanandum such that $\mathrm{KB}\models\varphi$, we define a monolithic explanation for $\varphi$ from $\mathrm{KB}$ as a minimal set of formulae that entails $\varphi$:
Definition 1.
(Monolithic Explanation) Let $\mathrm{KB}$ be a knowledge base and $\varphi$ an explanandum such that $\mathrm{KB}\models\varphi$. We say that $\epsilon\subseteq\mathrm{KB}$ is a monolithic explanation for $\varphi$ from $\mathrm{KB}$ if and only if: (i) $\epsilon\models\varphi$; and (ii) $\nexists\epsilon^{\prime}\subset\epsilon$ such that $\epsilon^{\prime}\models\varphi$.
Example 1.
Consider the knowledge base $\mathrm{KB}=\{p,\lnot p\vee q,\lnot p\vee r\}$ build up from $\mathcal{V}=\{p,q,r\}$. Notice that $\mathrm{KB}\models q$. Then, $\epsilon=\{p,\lnot p\vee q\}$ is a monolithic explanation for $q$ from $\mathrm{KB}$.
Note that in this paper we do not consider formulae $\epsilon\equiv\varphi$ as monolithic explanations. These trivial explanations, which are of the form “why $\varphi$, because $\varphi$, are uninformative pertaining the explanandum.
2.1.1 Duality of Minimal Unsatisfiable and Minimal Corrections Sets
Definition 2 (Minimal Unsatisfiable Set (MUS)).
Given an inconsistent knowledge base $\mathrm{KB}$, a subset $\mathcal{M}\subseteq\mathrm{KB}$ is an MUS if $\mathcal{M}$ is inconsistent and $\forall\mathcal{M}^{\prime}\subset\mathcal{M}$, $\mathcal{M}^{\prime}$ is consistent.
Definition 3 (Minimal Correction Set (MCS)).
Given an inconsistent knowledge base $\mathrm{KB}$, a subset $\mathcal{C}\subseteq\mathrm{KB}$ is an MCS if $\mathrm{KB}\setminus\mathcal{C}$ is consistent and $\forall\mathcal{C}^{\prime}\subset\mathcal{C}$, $\mathrm{KB}\setminus\mathcal{C}^{\prime}$ is inconsistent.
By definition, every inconsistent $\mathrm{KB}$ contains at least one MUS.
Definition 4 (Partial MUS).
A set of formulae $\Phi$ is a partial MUS of an inconsistent knowledge base $\mathrm{KB}$ if there exists at least one MUS $\mathcal{M}\subseteq\mathrm{KB}$ such that $\Phi\subseteq\mathcal{M}$.
Partial MUSes in an inconsistent knowledge base $\mathrm{KB}$ appear when a subset of formulae is set as hard, that is, formulae that must always be satisfied in a solution. Conversely, soft formulae may not always be satisfied. Given a formula $\varphi$, we will write $\varphi^{*}$ with $*\in\{s,h\}$ to denote it as soft and hard, respectively.
MUSes and MCSes are related by the concept of minimal hitting set:
Definition 5 (Minimal Hitting Set).
Given a collection $\Gamma$ of sets from a universe $U$, a hitting set for $\Gamma$ is a set $H\subseteq U$ such that $\forall S\in\Gamma$, $H\cap S\neq\emptyset$ and $\nexists H^{\prime}\subset H$ such that $H^{\prime}\cap S\neq\emptyset$.
The relationship between MUSes and MCSes is discussed by ? (?) and ? (?), and it was firstly presented by ? (?), where MUSes and MCSes are referred to as (minimal) conflicts and diagnoses, respectively.
Proposition 1.
A subset of an inconsistent knowledge base $\mathrm{KB}$ is an MUS (resp. MCS) if and only if it is a minimal hitting set of the collection of all MCSes (resp. MUSes) of $\mathrm{KB}$.
It follows from the above proposition that a cardinality minimal MUS (resp. MCS) is a minimal hitting set. Cardinality minimal MUS are referred to as SMUS, whereas a cardinality minimal MCS corresponds to the complement of a MaxSAT solution (?). We may refer to a cardinality minimal set as a minimum or smallest set.
Lemma 1.
Given a subset $\mathcal{H}$ of all the MCSes of knowledge base $\mathrm{KB}$, a hitting set is an SMUS if: (1) It is a minimal hitting set $h$ of $\mathcal{H}$, and (2) The subformula induced by $h$ is inconsistent.
See the work by ? (?) for a proof.
Proposition 1 and Lemma 1 naturally extend to the case of partial MUS. Note that when some formulae are set as hard in an inconsistent knowledge base, the set of all MCSes is a subset of the soft formulae. In this case, every minimal hitting set on the set of all MCSes is a partial MUS.
Finally, MUSes and monolithic explanations are related by the following:
Proposition 2.
Given a knowledge base $\mathrm{KB}$, a consistent set of formulae $\epsilon\subseteq\mathrm{KB}$ is a monolithic explanation for $\varphi$ from $\mathrm{KB}$ (Definition 1) if and only if $\epsilon$ is a partial MUS of $\epsilon\cup\{\lnot\varphi\}$.
Example 2.
Let $\mathrm{KB}=\{p,\lnot p\vee q,\lnot p\vee r\}$ and $\epsilon=\{p,\lnot p\vee q\}$ from Example 1. Notice how =$\mathcal{M}=\{p,\lnot p\vee q,\lnot q\}$ is an MUS of $\mathrm{KB}\cup\{\lnot q\}$. Then, it is easy to see that $\epsilon$ is a partial MUS of $\mathcal{M}$.
2.2 Modeling Uncertainty in Propositional Logic
Building on a propositional language $\mathcal{L}$, we can model the uncertainty of propositional formulae using a probability distribution over the possible worlds $\Omega$ of $\mathcal{L}$. Formally,
Definition 6 (Probability Distribution).
Let $\Omega$ be the set of possible worlds of the language $\mathcal{L}$. A probability distribution $P$ on $\Omega$ is a function $P:\Omega\mapsto[0,1]$ such that $\underset{\omega\in\Omega}{{{\sum}}}P(\omega)=1$.
In essence, a probability distribution over possible worlds creates a ranking between those worlds with respect to how likely they are to be true. This then allows us to quantify the uncertainty in a formula as follows:
Definition 7 (Degree of Belief).
Let $\Omega$ be the set of possible worlds and $P$ a probability distribution over $\Omega$. The degree of belief in a formula $\varphi\in\mathcal{L}$ is $P(\varphi)=\underset{\omega\models\varphi}{{{\sum}}}P(\omega)$.
We may refer to $P(\varphi)$ as degree of belief or probability of $\varphi$ interchangeably. Note that the possible worlds approach to probabilities is essentially equivalent to probabilities assigned directly to the formulae (?).
Now, the probability distribution on $\Omega$ can be induced from a weighted knowledge base, referred to as a belief base:
=$\mathcal{B}=\{(\phi_{1},w_{1}),\ldots,(\phi_{n},w_{n})\}$ (1)
where each formula $\phi_{i}\in\mathcal{L}$ is associated with a weight $w_{i}\in\mathbb{R}^{+}$. [^2] We assume, without loss of generality, that all weights are non-negative because a formula with a negative weight $w$ can be replaced by its negation with weight $-w$.
Intuitively, the weights serve as meta-information and reflect the certainty about the truth of the corresponding formulae – the higher the weight, the more certain the formula is. In that sense, formulae with higher weights are prioritized for satisfaction, effectively capturing the certainty of the particular formulae. This mechanism is especially useful for handling inconsistency and non-monotonic reasoning patterns, thus capturing a broader spectrum of problems. [^3] For example, the notion of inconsistency is relaxed as follows: Given two inconsistent formulae $\phi$ and $\lnot\phi$, if $P(\phi)=0.9$, then from the axioms of probability we have that $P(\lnot\phi)=0.1$. This then means that the worlds where $\lnot\phi$ is true are more unlikely than the worlds where $\phi$ is true, but not impossible. Further, we will denote with $\mathcal{B}^{\downarrow w}$ the classical projection of $\mathcal{B}$, that is, $\mathcal{B}^{\downarrow w}=\{\phi_{i}\>|\>(\phi_{i},w_{i})\in\mathcal{B}\}$.
Given a belief base $\mathcal{B}$, one way to induce a probability distribution is the following:
[…]
The induced probability distribution quantifies the likelihood that a given (possible) world is the actual world. Higher formula weights amplify the (log-) probability difference between a world that satisfies the formula and one that does not, other things being equal. Consequently, worlds that violate fewer formulas are deemed more probable. Note that a belief base $\mathcal{B}$ is essentially a log-linear model (?), from which a joint probability distribution of the set of variables of $\mathcal{L}$ is induced. Interestingly, log-linear models are special cases of Markov Logic Networks and can represent any positive distribution (?). When taken from context, we will simply use $P$ to denote the distribution induced from $\mathcal{B}$.
Entailment in a belief base $\mathrm{KB}$ becomes graded, that is, we now say that $\mathcal{B}$ entails a formula $\phi$ with degree of belief $P(\phi)$. However, when all weights are equal and tend to infinity, a belief base represents a uniform distribution over the worlds that satisfy it and, as such, entailment of a formula can be answered by computing the probability of the formula and checking whether it is $1$. In other words, entailment under belief bases collapses to classical entailment under knowledge bases. See ? (?) for a proof.
[…]
2.3 The Model Reconciliation Problem
The Model Reconciliation Problem (MRP), as introduced by ? (?), highlights the critical need for aligning the planning models of a human user and an agent to facilitate effective collaboration and understanding. This alignment becomes especially pertinent in scenarios where the agent’s plan deviates from human expectations, necessitating a mechanism to reconcile these differences through explanations. In this approach, the (planning) agent must have knowledge of the human’s model in order to contemplate their goals and foresee how its plan will be perceived by them. When there exist differences between the models of the agent and the human such that the agent’s plan diverges from the human’s expectations, the agent provides a minimal set of model differences, namely a model reconciling explanation, to the human.
It is important to highlight that, in order to effectively solve MRP, the following (implicit) assumptions typically hold:
  • 1.
    The agent model represents the ground truth or, in other words, the agent model is the “correct” encoding of the domain. This assumption is predicated on the notion that the explanation is generated from the agent’s perspective, thereby rendering it reasonable to assume that the agent “thinks” that its model is accurate or correct.
  • 2.
    The agent has access to the human model, which is an approximation of the actual human model. In the worst case, it can be empty; but, practically, it can be approximated based on past interactions (?, ?).
  • 3.
    Both models are assumed to be deterministic, and they thus are able to represent only deterministic domains. Note that while this is a restricting assumption, in Section 3.2 we present a framework that relaxes it.
Now, building upon the MRP foundation, ? (?) introduced its logic-based variant (L-MRP), where the models of the agent and the human user are represented as logical knowledge bases. As a model reconciling explanation must take into account both the knowledge base $\mathrm{KB}_{\alpha}$ of the agent providing an explanation as well as the knowledge base $\mathrm{KB}_{h}$ of the human receiving the explanation, it is defined slightly differently compared to monolithic explanations defined by Definition 1:
Definition 8 (Model Reconciling Explanation).
Given the knowledge bases of an agent $\mathrm{KB}_{\alpha}$ and a human user $\mathrm{KB}_{h}$ as well as an explanandum $\varphi$, such that $\mathrm{KB}_{\alpha}\models\varphi$ and $\mathrm{KB}_{h}\not\models\varphi$, $\mathcal{E}=\langle\epsilon^{+},\epsilon^{-}\rangle$ is a model reconciling explanation if and only if $\epsilon^{+}\subseteq\mathrm{KB}_{\alpha}$, $\epsilon^{-}\subseteq\mathrm{KB}_{h}$, and $(\mathrm{KB}_{h}\cup\epsilon^{+})\setminus\epsilon^{-}\models\varphi$.
When $\mathrm{KB}_{h}$ is updated with a model reconciling explanation $\mathcal{E}=\langle\epsilon^{+},\epsilon^{-}\rangle$, new formulae $\epsilon^{+}$ from $\mathrm{KB}_{\alpha}$ are added to $\mathrm{KB}_{h}$ and formulae $\epsilon^{-}$ from $\mathrm{KB}_{h}$ are retracted to ensure consistency. Note that since a model reconciling explanation is from the perspective of the agent’s knowledge base $\mathrm{KB}_{\alpha}$, we implicitly assume that if a formula in $\mathrm{KB}_{h}$ is inconsistent with $\mathrm{KB}_{\alpha}$, then that formula is “false” from the perspective of the agent.
Example 3.
Let $\mathrm{KB}_{\alpha}=\{a,\lnot a\vee b,\lnot a\vee c\}$ and $\mathrm{KB}_{h}=\{\lnot a,\lnot a\vee b\}$ be the knowledge bases of an agent and a human user, respectively, where $\mathrm{KB}_{\alpha}\models b$ and $\mathrm{KB}_{h}\not\models b$. A model reconciling explanation is then =$\mathcal{E}=\langle\{a\},\{\lnot a\}\rangle$, where $(\mathrm{KB}_{h}\cup\{a\})\setminus\{\lnot a\}=\{a,\lnot a\vee b\}\models b$.
3 A Framework for Probabilistic Explanation Generation
In this section, we outline a framework designed to extend the classical concepts of monolithic explanation, as defined by Definition 1, and model reconciling explanation, as defined by Definition 8, into probabilistic contexts.
3.1 Probabilistic Monolithic Explanations
Building on the classical notion of monolithic explanation presented in Definition 1, we introduce the concept of a probabilistic monolithic explanation. This concept aims to account for the uncertainty inherent in knowledge bases, providing a framework for explanations that not only identify contributing factors for an explanandum but also quantify the uncertainty in these factors. Throughout this section, we assume a belief base $\mathcal{B}$ and its induced probability distribution $P$.
Formally, a probabilistic monolithic explanation for an explanandum $\varphi$ from belief base $\mathcal{B}$ is defined as follows:
Definition 9 (Probabilistic Monolithic Explanation).
Let $\mathcal{B}$ be a belief base, $\mathcal{B}^{\downarrow w}$ its classical projection, and $\varphi$ an explanandum. We say that $\tilde{\epsilon}\subseteq\mathcal{B}^{\downarrow w}$ is a probabilistic monolithic explanation for $\varphi$ from $\mathcal{B}$ if and only if $P(\varphi\>|\>\tilde{\epsilon})>P(\varphi)$.
Intuitively, a probabilistic monolithic explanation $\tilde{\epsilon}$ seeks to increase the degree of belief in the explanandum $\varphi$. If $P(\varphi\>|\>\tilde{\epsilon})>P(\varphi)$, this then represents the case where $\tilde{\epsilon}$ increases the degree of belief in $\varphi$ and the greater the value of $P(\varphi\>|\>\tilde{\epsilon})$ the greater the degree of belief in $\varphi$.
Example 4.
Consider the belief base =$\mathcal{B}=\{(a,1),(\lnot a\vee b,2)\}$ and the explanandum $b$. The probability of the explanandum is $P(b)=0.73$. Then, $\tilde{\epsilon}_{1}=\{a\}$ and $\tilde{\epsilon}_{2}=\{\lnot a\vee b\}$ are two probabilistic monolithic explanations for $b$ from $\mathcal{B}$, that is, $P(b\>|\>\tilde{\epsilon}_{1})=0.88>P(b)$ and $P(b\>|\>\tilde{\epsilon}_{2})=0.78>P(b)$.
It is important to note that Definition 9 can be extended to the case where the formulae $\tilde{\epsilon}$ do not necessarily come from $\mathcal{B}$, but rather from the language $\mathcal{L}$. However, we restrict our attention only to formulae from $\mathcal{B}$ in order to be compatible with the classical notion of monolithic explanations (see Definition 1) and the algorithms that we will present in Section 4. For brevity, and until the end of this section, we will refer to probabilistic monolithic explanations as monolithic explanations.
Looking at Example 4, we can see that monolithic explanations will typically vary in their capacity to increase the degree of belief in the explanandum. In other words, each monolithic explanation provides us with some explanatory gain for the explanandum. Following ? (?, ?), explanatory gain is defined as follows:555 ? (?) originally introduced this measure to quantify the (weak) explanatory power of a hypothesis with respect to evidence, essentially evaluating how effectively the hypothesis explains the evidence.
Definition 10 (Explanatory Gain of Monolithic Explanations).
Let $\tilde{\epsilon}$ be a monolithic explanation for explanandum $\varphi$ from belief base $\mathcal{B}$. The explanatory gain of $\tilde{\epsilon}$ for $\varphi$ is defined as $G(\tilde{\epsilon},\varphi)=\log\left(\frac{P(\varphi\>|\>\tilde{\epsilon})}{P(\varphi)}\right)$. [^6] We use $\log$ with base $2$ in our calculations., [^7] Note that $G(\tilde{\epsilon},\varphi)$ is always positive due to the requirement of monolithic explanations that $P(\varphi\>|\>\tilde{\epsilon})>P(\varphi)$ (Definition 9).
In essence, the explanatory gain can be thought of as a measure that quantifies how well the monolithic explanation $\tilde{\epsilon}$ explains the explanandum $\varphi$ or, equivalently, the degree to which $\tilde{\epsilon}$ entails $\varphi$. The greater the value of $G(\tilde{\epsilon},\varphi)$, the more substantial the explanatory gain and, hence, the more effective $\tilde{\epsilon}$ is at explaining $\varphi$.
It is essential to recognize that this measure, while initially introduced to assess the weak explanatory power of hypotheses in light of evidence (?), it is used here to evaluate monolithic explanations. By quantifying the extent to which a monolithic explanation explains an explanandum, we can systematically identify the most informative monolithic explanations within a probabilistic framework.
Example 5.
Continuing from Example 4, consider the monolithic explanations $\tilde{\epsilon}_{1}=\{a\}$, $\tilde{\epsilon}_{2}=\{\lnot a\vee b\}$, and $\tilde{\epsilon}_{3}=\{a,\lnot a\vee b\}$ for explanandum $b$. The explanatory gains of $\tilde{\epsilon}_{1}$, $\tilde{\epsilon}_{2}$, and $\tilde{\epsilon}_{3}$ for $b$ are $G(\tilde{\epsilon}_{1},b)=\log\left(\frac{P(b\>|\>\tilde{\epsilon}_{1})}{P(b)}\right)=\log\left(\frac{0.88}{0.73}\right)=0.27$, $G(\tilde{\epsilon}_{2},b)=\log\left(\frac{P(b\>|\>\tilde{\epsilon}_{2})}{P(b)}\right)=\log\left(\frac{0.78}{0.73}\right)=0.11$, and $G(\tilde{\epsilon}_{3},b)=\log\left(\frac{P(b\>|\>\tilde{\epsilon}_{3})}{P(b)}\right)=\log\left(\frac{1}{0.73}\right)=0.45$, respectively.
Now, a natural course of action when seeking monolithic explanations for an explanandum is to seek the one with the highest explanatory gain. While it is tempting to do this, it is important to emphasize that when a monolithic explanation entails the explanandum, then the explanatory gain takes on its greatest value. For example,
Example 6.
Consider the three monolithic explanations $\tilde{\epsilon}_{1}$, $\tilde{\epsilon}_{2}$, and $\tilde{\epsilon}_{3}$ from Example 5. Notice that $\tilde{\epsilon}_{3}=\{a,\lnot a\vee b\}$ entails $b$ (i.e., $\tilde{\epsilon}_{3}\models b$) and that its explanatory gain is higher than that of $\tilde{\epsilon}_{1}$ and $\tilde{\epsilon}_{2}$. As $\tilde{\epsilon}_{1}$, $\tilde{\epsilon}_{2}$, and $\tilde{\epsilon}_{3}$ are the only three possible explanations for $b$, $G(\tilde{\epsilon}_{3},b)$ is indeed the maximum achievable explanatory gain for $b$.
We formalize this in the following proposition:
Proposition 3.
Given a monolithic explanation $\tilde{\epsilon}$ for an explanandum $\varphi$ from belief base $\mathcal{B}$, if $\tilde{\epsilon}\models\varphi$, then $G(\tilde{\epsilon},\varphi)$ achieves its maximal value for $\varphi$, specifically $G(\tilde{\epsilon},\varphi)=-\log P(\varphi)$.
Proof.
If $\tilde{\epsilon}\models\varphi$, then for all possible worlds $\omega$ in which $\omega\models\tilde{\epsilon}$, it holds that $\omega\models\varphi$. That is, the worlds $\omega$ in which $\tilde{\epsilon}$ is true are subsumed by the worlds in which $\varphi$ is true, which implies that also $\omega\models\varphi\land\tilde{\epsilon}$. Consequently, $P(\varphi\>|\>\tilde{\epsilon})=\frac{\underset{\omega\models\varphi\land\tilde{\epsilon}}{{{\sum}}}P(\omega)}{\underset{\omega\models\tilde{\epsilon}}{{{\sum}}}P(\omega)}=\frac{\underset{\omega\models\tilde{\epsilon}}{{{\sum}}}P(\omega)}{\underset{\omega\models\tilde{\epsilon}}{{{\sum}}}P(\omega)}=1$. Therefore, when $\tilde{\epsilon}\models\varphi$, the explanatory gain of $\tilde{\epsilon}$ for $\varphi$ is $G(\tilde{\epsilon},\varphi)=\log\left(\frac{P(\varphi\>|\>\tilde{\epsilon})}{P(\varphi)}\right)=\log\left(\frac{1}{P(\varphi)}\right)=-\log P(\varphi)$. ∎
The following corollary follows naturally from Proposition 3:
Corollary 1.
Let $\tilde{E}(\varphi)$ denote the set of all monolithic explanations for explanandum $\varphi$ from belief base $\mathcal{B}$. For any two monolithic explanations $\tilde{\epsilon}_{1},\tilde{\epsilon}_{2}\in\tilde{E}(\varphi)$, if $\tilde{\epsilon}_{1}\models\varphi$ and $\tilde{\epsilon}_{2}\models\varphi$ (resp. $\tilde{\epsilon}_{2}\not\models\varphi$), then $G(\tilde{\epsilon}_{1},\varphi)=G(\tilde{\epsilon}_{2},\varphi)$ (resp. $G(\tilde{\epsilon}_{1},\varphi)>G(\tilde{\epsilon}_{2},\varphi)$).
What Proposition 3 and Corollary 1 essentially underscore is that the exclusive focus on explanatory gain as an evaluation metric for a monolithic explanation neglects the inherent likelihood of the explanation itself. That is, the explanatory gain of a monolithic explanation for an explanandum evaluates how effectively the explanation explains the explanandum, assuming that the explanation itself is true. Nonetheless, this premise often lacks practical relevance because, in probabilistic contexts, each monolithic explanation is associated with a probability reflecting its likelihood for being true. Therefore, a good measure for evaluating monolithic explanations should incorporate the explanation’s inherent plausibility.
Addressing this gap, ? (?) introduced the concept of (strong) explanatory power that integrates the monolithic explanation’s explanatory gain with its probability, offering a more balanced metric for evaluating monolithic explanations. [^8] Good’s measure of (strong) explanatory power is defined as ${{\log}}[\frac{P(\varphi\>|\>h)\cdot P(h)^{\gamma}}{P(\varphi)}]$, where $h$ is a hypothesis and $10<\gamma<$ 1 a constant (?). Building on Good’s measure of explanatory power, we adapt it to our setting and define it as follows: [^9] For a detailed defense of Good’s measure as a quantitative criterion for explanatory power, alongside a discussion of relevant properties and a comprehensive comparison with other measures, we refer the reader to the work by ? (?).
Definition 11 (Explanatory Power of Monolithic Explanations).
Let $\tilde{\epsilon}$ be a monolithic explanation for explanandum $\varphi$ from belief base $\mathcal{B}$. The explanatory power of $\tilde{\epsilon}$ for $\varphi$ is defined as $\mathrm{EP}(\tilde{\epsilon},\varphi)=G(\tilde{\epsilon},\varphi)+\gamma\cdot P(\tilde{\epsilon})$, where $\gamma\in[0,1]$ is a constant.
This definition effectively combines the measure of how much a monolithic explanation explains the explanandum (explanatory gain) with the likelihood of the explanation itself, mediated by a parameter $\gamma$. The constant $\gamma$ serves as a tuning parameter, enabling the adjustment of the relative importance of the monolithic explanation’s probability in the overall assessment of explanatory power. This flexibility is important for tailoring the evaluation process to specific contexts or preferences, where the balance between the informativeness of a monolithic explanation and its plausibility may vary.
Example 7.
Consider the belief base =$\mathcal{B}=\{(a,1.5),(b,3),(\lnot a\vee c,1),(\lnot b\vee c,1)\}$ and the explanandum $c$ with initial probability $P(c)=0.84$. Notice that $\tilde{\epsilon}_{1}=\{a,\lnot a\vee c\}$ and $\tilde{\epsilon}_{2}=\{b,\lnot b\vee c\}$ are two monolithic explanations for $c$ from $\mathcal{B}$, each of which entail $c$ (i.e., $\tilde{\epsilon}_{1}\models c$ and $\tilde{\epsilon}_{2}\models c$), with probabilities $P(\tilde{\epsilon}_{1})=0.68$ and $P(\tilde{\epsilon}_{2})=0.80$, respectively. This means that their explanatory gain for $c$ is equal (Corollary 1), that is, $G(\tilde{\epsilon}_{1},c)=G(\tilde{\epsilon}_{2},c)=0.25$. Now, assuming $\gamma=0.5$, the explanatory power of $\tilde{\epsilon}_{1}$ and $\tilde{\epsilon}_{2}$ respectively is $\mathrm{EP}(\tilde{\epsilon}_{1},c)=0.25+0.5\cdot 0.68=0.59$ and $\mathrm{EP}(\tilde{\epsilon}_{2},c)=0.25+0.5\cdot 0.80=0.65$.
With the introduction of explanatory power as an evaluative measure of (probabilistic) monolithic explanations, we can now define a (probabilistic) preference relation among monolithic explanations, which allows for a systematic approach to determining the most effective monolithic explanation for a given explanandum:
Definition 12 (Preference Relation for Monolithic Explanation).
Let $\tilde{\epsilon}_{1}$ and $\tilde{\epsilon}_{2}$ be two monolithic explanations for explanandum $\varphi$ from belief base $\mathcal{B}$. $\tilde{\epsilon}_{1}$ is preferred over $\tilde{\epsilon}_{2}$, denoted as $\tilde{\epsilon}_{1}\succeq\tilde{\epsilon}_{2}$, if and only if $\mathrm{EP}(\tilde{\epsilon}_{1},\varphi)\geq\mathrm{EP}(\tilde{\epsilon}_{2},\varphi)$.
This definition enables a quantitatively grounded approach to preference among monolithic explanations, where the preference is directly tied to the explanatory power of each explanation. It facilitates a structured way to navigate the space of potential monolithic explanations, prioritizing those that not only explain the explanandum more effectively, but also align more closely with the existing knowledge represented by the belief base $\mathcal{B}$.
Example 8.
Continuing from Example 7, the two monolithic explanations for $c$ from $\mathcal{B}$ are $\tilde{\epsilon}_{1}$ and $\tilde{\epsilon}_{2}$ and have explanatory power $\mathrm{EP}(\tilde{\epsilon}_{1},c)=0.59$ and $\mathrm{EP}(\tilde{\epsilon}_{2},c)=0.65$. Thus, $\tilde{\epsilon}_{2}$ is preferred over $\tilde{\epsilon}_{1}$ (i.e., $\tilde{\epsilon}_{2}\succeq\tilde{\epsilon}_{1}$).
Finally, given the set of all monolithic explanations for an explanandum, we say that a monolithic explanation is most preferred if and only if it is (probabilistically) preferred over every other possible monolithic explanation for that explanandum. Formally,
Definition 13 (Most-Preferred Monolithic Explanation).
Let $\tilde{E}(\varphi)$ denote the set of all monolithic explanations for explanandum $\varphi$ from belief base $\mathcal{B}$. A monolithic explanation $\tilde{\epsilon}^{*}\in\tilde{E}(\varphi)$ is the most-preferred monolithic explanation if and only if $\tilde{\epsilon}^{*}\succeq\tilde{\epsilon}$ for all $\tilde{\epsilon}\in\tilde{E}(\varphi)$.
In the next section, we consider how probabilistic monolithic explanations will look like for the model reconciliation problem.
3.2 Probabilistic Model Reconciling Explanations
Recall from Section 2.3 that, in the model reconciliation problem (MRP), the models of the agent and the human user diverge with respect to an explanandum, insofar as the explanandum is explicable in the agent’s model but inexplicable in the human’s model. The goal is then to find a model reconciling explanation (i.e., a set of model differences) such that the explanandum becomes explicable in the human’s model. Three important assumptions underlying MRP typically hold: (1) the agent model is the ground truth; (2) the agent has access to the human model; and (3) both models are deterministic.
As we described in Section 2.3, assumption (1) is reasonable since explanations are generated from the agent’s perspective. In other words, the agent “thinks” that its model is correct. For assumption (2), the agent does not have access to the human’s actual model, but an approximation of it. In the worst case, it can be empty; but, practically, it can be approximated based on past interactions (?, ?). For assumption (3), we will relax the assumption that the human model is deterministic in our work, but we will still assume that the agent model is deterministic.
The motivation for moving away from deterministic human models becomes stronger when we consider two key points. First, since the agent is using an approximated human model, deterministic approximations are more likely to be inaccurate compared to probabilistic ones. Consequently, deterministic models may generate explanations that are incorrect or not meaningful for the user, thereby reducing the effectiveness of MRP. Secondly, it is likely that humans hold beliefs with varying degrees of certainty, highlighting a shortfall of deterministic models in capturing this range of uncertainties. These factors together underscore the necessity for models that incorporate probabilistic aspects, thus potentially enabling a more accurate and user-relevant application of MRP.
To that end, we will now expand the scope of MRP to cases in which the agent is uncertain about the human model. Particularly, we build on our previous work (?), wherein both the agent and human models are represented as (logical) knowledge bases (see Definition 8), and extend it to the case where the human knowledge base is probabilistic (i.e., a belief base). In other words, we are now interested in probabilistic model reconciling explanations.
First, we show through the following example how the concepts surrounding probabilistic monolithic explanations introduced in the previous section are applicable to the case of an agent knowledge base $\mathrm{KB}_{\alpha}$ and a human belief base $\mathcal{B}_{h}$.
Example 9.
Let $\mathrm{KB}_{\alpha}=\{a,\lnot a\vee b,c\}$ and $\mathcal{B}_{h}=\{(c,2),(\lnot c\vee\lnot a,2)\}$ be the knowledge bases of an agent and the belief base of a human, respectively. Additionally, let $b$ be the explanandum, where $\mathrm{KB}_{\alpha}\models b$ and $P_{h}(b)=0.5$. The goal in this example would then be to find which formulae from $\mathrm{KB}_{\alpha}$ increase the probability of the explanandum for $\mathcal{B}_{h}$, that is, to find a probabilistic monolithic explanation $\tilde{\epsilon}$ for $b$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$ such that $P_{h}(b\>|\>\tilde{\epsilon})>P_{h}(b)$ (Definition 9).
Given $\mathrm{KB}_{\alpha}$, there are three possible monolithic explanations: $\tilde{\epsilon}_{1}=\{a\}$, $\tilde{\epsilon}_{2}=\{\lnot a\vee b\}$, and $\epsilon_{3}=\{a,\lnot a\vee b\}$. Evaluating them with respect to the probability distribution induced by $\mathcal{B}_{h}$, we get $P_{h}(b\>|\>\tilde{\epsilon}_{1})=0.5$, $P_{h}(b\>|\>\tilde{\epsilon}_{2})=0.55$, and $P_{h}(b\>|\>\tilde{\epsilon}_{3})=1$. Notice now that only $\tilde{\epsilon}_{2}$ and $\tilde{\epsilon}_{3}$ qualify as monolithic explanations since $P_{h}(b\>|\>\tilde{\epsilon}_{2})>P_{h}(b)$ and $P_{h}(b\>|\>\tilde{\epsilon}_{3})>P_{h}(b)$, whilst $\tilde{\epsilon}_{1}$ does not qualify as a monolithic explanation as $P_{h}(b\>|\>\tilde{\epsilon}_{1})=P_{h}(b)=0.5$.
Given $\tilde{\epsilon}_{2}$ and $\tilde{\epsilon}_{3}$ as the two possible monolithic explanations, we can now evaluate their effectiveness in terms of explanatory gain (Definition 10) and explanatory power (Definition 11). In terms of explanatory gain, we get $G(\tilde{\epsilon}_{2},b)=0.14$ and $G(\tilde{\epsilon}_{3},b)=1$. In terms of explanatory power (for $\gamma=0.5$), we get $\mathrm{EP}(\tilde{\epsilon}_{2},b)=0.59$ and $\mathrm{EP}(\tilde{\epsilon}_{3},b)=1.04$. Finally, following the definition of most-preferred monolithic explanation (Definition 13), we get that $\tilde{\epsilon}_{3}$ is the most-preferred monolithic explanation for $b$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$.
On the one hand, example 9 shows that the definitions introduced in Section 3.1 can be directly applied to the case of an agent knowledge base $\mathrm{KB}_{\alpha}$ and a human belief base $\mathcal{B}_{h}$. On the other hand, there is something important to highlight here. Despite $\tilde{\epsilon}_{3}$ being the most-preferred monolithic explanation (i.e., it has the highest explanatory power), notice that its probability $P_{h}(\tilde{\epsilon}_{3})=0.09$ is rather low, which means that its negation $\lnot\tilde{\epsilon}_{3}$ has a much higher probability with $P_{h}(\lnot\tilde{\epsilon}_{3})=0.91$. Logically, this is explained by the fact that $\tilde{\epsilon}_{3}$ is inconsistent with the formulae in $\mathcal{B}_{h}^{\downarrow w}$. Therefore, the probabilistic monolithic explanation $\tilde{\epsilon}_{3}$ may not achieve the intended “reconciliation” between the agent and the human.
Recall that a model reconciling explanation (see Definition 8) is of the form $\mathcal{E}=\langle\epsilon^{+},\epsilon^{-}\rangle$, where $\epsilon^{-}$ is specifically intended to resolve the inconsistency between the agent and the human with respect to the explanandum. Intuitively, the provision of $\epsilon^{-}$ can be thought of as the agent’s suggestion of what is “false” in the human knowledge base, at least compared to the agent knowledge base. In the case of a human belief base $\mathcal{B}_{h}$, we can account for $\epsilon^{-}$ by finding a set of formulae from $\mathcal{B}_{h}$ such that $P_{h}(\epsilon^{+}\>|\>\lnot\epsilon^{-})>P_{h}(\epsilon^{+})$. For example,
Example 10.
Let $\mathrm{KB}_{\alpha}=\{a,\lnot a\vee b,c\}$ and $\mathcal{B}_{h}=\{(c,2),(\lnot c\vee\lnot a,2)\}$ from Example 9. From the perspective of $\mathrm{KB}_{\alpha}$, explanation $\tilde{\epsilon}=\{a,\lnot a\vee c\}$ can be seen as the formulae that should be true (e.g., added) in $\mathcal{B}_{h}$ (i.e., $\tilde{\epsilon}^{+}=\tilde{\epsilon}$). However, notice that $\tilde{\epsilon}^{+}$ is inconsistent with $\mathcal{B}_{h}^{\downarrow w}=\{c,\lnot c\vee\lnot a\}$. Thus, from the perspective of $\mathrm{KB}_{\alpha}$, some formulae from $\mathcal{B}_{h}^{\downarrow w}$ are false (e.g., they should be retracted). One can see that $\tilde{\epsilon}^{-}=\lnot c\vee\lnot a$ is the only formula that should be false as it is the only one that is inconsistent with $\mathrm{KB}_{\alpha}$. Indeed, if $\tilde{\epsilon}^{-}$ is assumed to be false, then the probability of $\tilde{\epsilon}^{+}$ increases, i.e., $P_{h}(\tilde{\epsilon}^{+}\>|\>\lnot\tilde{\epsilon}^{-})=0.5>P_{h}(\tilde{\epsilon}^{+})=0.09$. Therefore, $\tilde{\epsilon}^{+}$ and $\tilde{\epsilon}^{-}$ can be seen as a model reconciling explanation for $b$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$.
Before formally defining what constitutes a probabilistic model reconciling explanation, we state the following assumptions underlying our framework:
  • $\bullet$
    Shared Domain Language: The agent and the human user share the same (propositional) language $\mathcal{L}$, that is, they share the same set of atomic variables $\mathcal{V}$ from which formulae specific to a domain can be constructed.
  • $\bullet$
    Agent Knowledge Base: The agent model is represented by the (deterministic) knowledge base $\mathrm{KB}_{\alpha}$, encoding the ground truth of the domain.
  • $\bullet$
    Human Belief Base: The human model is represented by the belief base $\mathcal{B}_{h}$ (and its associated probability distribution $P_{h}$), reflecting the agent’s uncertainty, for example, its degrees of belief about the human model. The agent has access to $\mathcal{B}_{h}$ a-priori. [^10] We leave the question of acquiring (or learning) the human belief base open for future work.
We define a probabilistic model reconciling explanation as follows:
Definition 14 (Probabilistic Model Reconciling Explanation).
Given the knowledge base $\mathrm{KB}_{\alpha}$ of an agent, the belief base $\mathcal{B}_{h}$ of a human user, and an explanandum $\varphi$ such that $\mathrm{KB}_{\alpha}\models\varphi$ and $P_{h}(\varphi)<1$, ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ is a probabilistic model reconciling explanation if and only if $\tilde{\epsilon}^{+}\subseteq\mathrm{KB}_{\alpha}$ and $\tilde{\epsilon}^{-}\subseteq\mathcal{B}_{h}^{\downarrow w}$, and $P_{h}(\varphi\>|\>\tilde{\epsilon}^{+})>P_{h}(\varphi)$ and $P_{h}(\tilde{\epsilon}^{+}\>|\>\lnot\tilde{\epsilon}^{-})>P_{h}(\tilde{\epsilon}^{+})$.
A probabilistic model reconciling explanation ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$ is a tuple that increases the degree of belief in $\varphi$ with $\tilde{\epsilon}^{+}$, as well as increasing the degree of belief in $\tilde{\epsilon}^{+}$ with $\tilde{\epsilon}^{-}$ if $\tilde{\epsilon}^{+}$ is inconsistent with $\mathcal{B}_{h}^{\downarrow w}$. For brevity, until the end of this section, we will refer to probabilistic model reconciling explanations $\tilde{\mathcal{E}}$ as model reconciling explanations. In this context, the notion of explanatory gain takes the following form:
Definition 15 (Explanatory Gain for Model Reconciling Explanations).
Let ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ be a model reconciling explanation for explanandum $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$. The explanatory gain of $\tilde{\mathcal{E}}$ for $\varphi$ is defined as $G(\tilde{\mathcal{E}},\varphi)=\log\left(\frac{P(\varphi\>|\>\tilde{\epsilon}^{+})}{P(\varphi)}\right)+\log\left(\frac{P(\tilde{\epsilon}^{+}\>|\>\lnot\tilde{\epsilon}^{-})}{P(\tilde{\epsilon}^{+})}\right)$.
In essence, the explanatory gain of ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ evaluates to what extent $\tilde{\epsilon}^{+}$ increases the probability of $\varphi$, as well as the extent to which $\tilde{\epsilon}^{-}$ increases the probability of $\tilde{\epsilon}^{+}$, assuming that $\tilde{\epsilon}^{-}$ is false.
Example 11.
Let ~=$\tilde{\mathcal{E}}=\langle\{a,\lnot a\vee b\},\{\lnot c\vee\lnot a\}\rangle$ be the model reconciling explanation for $b$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$ in Example 10. The explanatory gain of $\tilde{\mathcal{E}}$ for $b$ is $G(\tilde{\mathcal{E}},b)=\log\left(\frac{1}{0.5}\right)+\log\left(\frac{0.5}{0.09}\right)=1+2.47=3.47$.
Similarly, the notion of explanatory power is defined in the following way:
Definition 16 (Explanatory Power for Model Reconciling Explanations).
Let ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ be a model reconciling explanation for explanandum $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$. The explanatory power of $\tilde{\mathcal{E}}$ for $\varphi$ is defined as $\mathrm{EP}(\tilde{\mathcal{E}},\varphi)=G_{h}(\tilde{\mathcal{E}},\varphi)+\gamma\cdot\left(P_{h}(\tilde{\epsilon}^{+})+P_{h}(\tilde{\epsilon}^{-})\right)$, where $\gamma\in[0,1]$ is a constant.
This definition of explanatory power of ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ assesses, in addition to the explanatory gain of $\tilde{\mathcal{E}}$, the likelihoods of $\tilde{\epsilon}^{+}$ and $\tilde{\epsilon}^{-}$, with $\gamma$ parameterizing their relative importance in the overall assessment.
Example 12.
Continuing from Example 11, the explanatory power of ~=$\tilde{\mathcal{E}}=\langle\{a,\lnot a\vee b\},\{\lnot c\vee\lnot a\}\rangle$ for $b$ (for $\gamma=0.5$) is $\mathrm{EP}(\tilde{\mathcal{E}},b)=3.47+0.5\cdot(0.09+0.90)=3.96$
Finally, a preference relation and a most-preferred model reconciling explanation can be defined in the same manner as in Definition 12 and Definition 13, respectively.
Definition 17 (Preference Relation for Model Reconciling Explanation).
Let ~$\tilde{\mathcal{E}}_{1}$ and ~$\tilde{\mathcal{E}}_{2}$ be two model reconciling explanations for explanandum $\varphi$ from knowledge base $\mathrm{KB}_{\alpha}$ for beleif base $\mathcal{B}_{h}$. ~$\tilde{\mathcal{E}}_{1}$ is preferred over ~$\tilde{\mathcal{E}}_{2}$, denoted ~$\tilde{\mathcal{E}}_{1}\succeq\tilde{\mathcal{E}}_{2}$, if and only if $\mathrm{EP}(\tilde{\mathcal{E}}_{1})\geq\mathrm{EP}(\tilde{\mathcal{E}}_{2})$.
Definition 18 (Most-Preferred Model Reconciling Explanation).
Let $\tilde{E}(\varphi)$ denote the set of all model reconciling explanations for explanandum $\varphi$ from knowledge base $\mathrm{KB}_{\alpha}$ for belief base $\mathcal{B}_{h}$. A model reconciling explanation ~$\tilde{\mathcal{E}}^{*}\in\tilde{E}(\varphi)$ is the most-preferred model reconciling explanation for $\varphi$ if and only if $\tilde{\mathcal{E}}^{*}\succeq\tilde{\mathcal{E}}$ for all ~$\tilde{\mathcal{E}}^{\in}\tilde{E}(\varphi)$.
4 Computing Explanations
We now describe algorithms for computing explanations. We first review two algorithms proposed in our previous work (?) for computing classical (deterministic) monolithic explanations (Definition 1) and model reconciling explanations (Definition 8), and then show how to extend them to the probabilistic case.
4.1 Classical Explanations
We previously introduced an approach for computing minimum size monolithic explanations for an explanandum $\varphi$ from a knowledge base $\mathrm{KB}$, where $\mathrm{KB}\models\varphi$ (?). The principal idea of this approach is to reduce the problem of computing a monolithic explanation of minimum size to the one of computing a smallest minimal unsatisfiable set (SMUS) over an inconsistent knowledge base.
In particular, notice that, by definition, we have that $\mathrm{KB}\models\varphi$ if and only if $\mathrm{KB}\cup\{\lnot\varphi\}$ is inconsistent. Moreover, in Proposition 2, we have already stated the relation between a monolithic explanation and a minimal unsatisfiable set (MUS). This suggests that, in order to extract a monolithic explanation, we just need to run an MUS solver over the knowledge base $\mathrm{KB}^{s}\cup\{\lnot\varphi^{h}\}$, where $\mathrm{KB}^{s}$ and $\varphi^{h}$ denote that $\mathrm{KB}$ and $\varphi$ are treated as soft and hard constraints, respectively, and then remove $\lnot\varphi$ from the returned MUS. [^11] Recall that soft constraints may be removed by the MUS solver, while hard constraints will not be removed. The hitting set duality relating MUSes and minimal correction sets (MCSes) (see Lemma 1) is a key aspect for the computation of an SMUS.
Input: Knowledge base $\mathrm{KB}$ and explanandum $\varphi$
Result: A minimum size monolithic explanation $\epsilon$ for $\varphi$ from $\mathrm{KB}$
1
$\mathcal{H}\leftarrow\emptyset$
3 while true do
       $seed\leftarrow\texttt{minHS}(\mathcal{H})$
        // compute a minimal hitting set
4       $\epsilon\leftarrow\{c_{i}\;|\;i\in seed\}$
5       if not SAT$(\epsilon\cup\{\lnot\varphi\})$ then
             return $\epsilon$
              // minimum size monolithic explanation
6
7       else
             $\mathcal{C}\leftarrow\texttt{getMCS}(seed,\mathrm{KB}^{s}\cup\{\lnot\varphi^{h}\})$
              // compute a minimal correction set
8
9      $\mathcal{H}\leftarrow\mathcal{H}\cup\{\mathcal{C}\}$
10
Algorithm 1 monolithic-explanation$(\mathrm{KB},\varphi)$
Table 1: Example of Algorithm 1 for computing a monolithic explanation of minimum size.
$\mathrm{KB}=\{\;\stackrel{{\scriptstyle C_{1}}}{{a\lor b}},\stackrel{{\scriptstyle C_{2}}}{{\lnot b\lor c}},\stackrel{{\scriptstyle C_{3}}}{{\lnot c}},\stackrel{{\scriptstyle C_{4}}}{{\lnot b\lor d}}\}$
$\mathrm{KB}\models a$
1. $\mathcal{H}\leftarrow\emptyset$
2. $seed\leftarrow\emptyset$ $minHS(\mathcal{H})$
3. ⊧̸$\emptyset\not\models a∅ ⊧$ $SAT(\epsilon\cup\{\lnot a\})$
4. $\mathcal{C}\leftarrow\{C_{1}\}$ MCS computed on $KB^{s}\cup\{\lnot a^{h}\}$ starting with the seed $seed$
5. $\mathcal{H}\leftarrow\{\{C_{1}\}\}$
6. $seed\leftarrow\{C_{1}\}$ $minHS(\mathcal{H})$
7. $\{a\vee b\}\not\models a$ $SAT(\epsilon\cup\{\lnot a\})$
8. $\mathcal{C}\leftarrow\{C_{2}\}$ MCS computed on $KB^{s}\cup\{\lnot a^{h}\}$ starting with the seed $seed$
9. $\mathcal{H}\leftarrow\{\{C_{1}\},\{C_{2}\}\}$
10. $seed\leftarrow\{C_{1},C_{2}\}$ $minHS(\mathcal{H})$
11. $\{a\vee b,\lnot b\vee c\}\not\models a$ $SAT(\epsilon\cup\{\lnot a\})$
12. $\mathcal{C}\leftarrow\{C_{3}\}$ MCS computed on $KB^{s}\cup\{\lnot a^{h}\}$ starting with the seed $seed$
13. $\mathcal{H}\leftarrow\{\{C_{1}\},\{C_{2}\},\{C_{3}\}\}$
14. $seed\leftarrow\{C_{1},C_{2},C_{3}\}$ $minHS(\mathcal{H})$
15. $\{a\vee b,\lnot b\vee c,\lnot c\}\models a$$\lnot SAT(\epsilon\cup\{\lnot a\})$
16. $Return$$\{a\vee b,\lnot b\vee c,\lnot c\}$ minimum size monolithic explanation for $a$ from $\mathrm{KB}$
Algorithm 1 describes the main steps of our approach. $\mathcal{H}$ is a collection of sets, where each set corresponds to an MCS on $\mathrm{KB}$. At the beginning, it is initialized with the empty set (line 1). Each MCS in $\mathcal{H}$ is represented as the set of the indexes of the formulae in it. $\mathcal{H}$ stores the MCSes computed so far. At each step, a minimal hitting set on $\mathcal{H}$ is computed (line 1). In line 1, the formulae induced by the computed minimal hitting set is stored in $\epsilon$. Then, $\epsilon\cup\{\lnot\varphi\}$ is evaluated for satisfiability (line 1). If $\epsilon\cup\{\lnot\varphi\}$ is inconsistent, then $\epsilon$ is a monolithic explanation of minimum size and the algorithm returns $\epsilon$. If instead $\epsilon\cup\{\lnot\varphi\}$ is consistent, then it means that $\epsilon\not\models\varphi$ and the algorithm continues in line 1. The computation of an MCS of this kind can be performed via standard MCS procedures (?), using the set of formulae indexed by the seed as the starting formula to extend. Since $\varphi$ is set to hard (line 1), the returned MCS $\mathcal{C}$ is guaranteed to be contained in $\mathrm{KB}$. Due to the hitting set duality relation, we will also have $\epsilon\subseteq\mathrm{KB}$. Finally, notice that the procedure $getMCS$ always reports a new MCS because, by construction, we have $seed\subseteq\mathrm{KB}\setminus\mathcal{C}$. In fact, the $seed$ contains at least one formula for each previously computed MCS and, thus, $seed\cap\mathcal{C}=\emptyset$ (i.e., at least one formula for each previously computed MCS is not in $\mathcal{C}$).
Algorithm 1 is complete in the sense that eventually a monolithic explanation $\epsilon\subseteq\mathrm{KB}$ of minimum size such that $\epsilon\models\varphi$ will be returned. This can be easily verified by observing that every time $\epsilon\cup\{\lnot\varphi\}$ is satisfiable, a new MCS is computed. Eventually, all the MCSes will be computed and, from Propositions 1 and 2, it follows that a minimal hitting set on the collection of all MCSes corresponds to the smallest MUS, and as such, to a monolithic explanation of minimum size.
Note that deciding whether there exists a monolithic explanation of size less or equal to $k$ is $\Sigma_{2}^{p}$-complete and extracting a smallest monolithic explanation is in $FP^{\Sigma_{2}^{p}}$. This follows directly from the complexity of deciding and computing an SMUS on which Algorithm 1 is based on (?).
Input: Knowledge bases $\mathrm{KB}_{\alpha}$ and $\mathrm{KB}_{h}$ and explanandum $\varphi$
Result: A model reconciling explanation $\mathcal{E}=\langle\epsilon^{+},\epsilon^{-}\rangle$ for $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathrm{KB}_{h}$
1 $\mathbb{R}\leftarrow\emptyset$
2 $\mathrm{KB}_{\alpha}^{h}\leftarrow\mathrm{KB}_{\alpha}\cap\mathrm{KB}_{h}$
3 $\mathrm{KB}_{\alpha}^{s}\leftarrow\mathrm{KB}_{\alpha}\setminus\mathrm{KB}_{\alpha}^{h}$
4 if not SAT$(\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha})$ then
       $E^{-}\leftarrow\texttt{getMCS}((\mathrm{KB}_{h}\setminus\mathrm{KB}_{\alpha})^{s}\cup\mathrm{KB}_{\alpha}^{h})$
        // restore consistency on $\mathrm{KB}_{h}$
5       $\mathrm{KB}_{h}\leftarrow\mathrm{KB}_{h}\setminus E^{-}$
6
7while true do
8       $seed\leftarrow\texttt{minHS}(\mathbb{R})$
       $\epsilon^{+}\leftarrow\{c_{i}\;|\;i\in seed\}$
        // explanation $\epsilon^{+}$ induced by the seed
9       if not SAT$(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot\varphi\})$ then
10             $\epsilon^{-}\leftarrow\emptyset$
11             if not SAT$(\mathrm{KB}_{h}\cup\epsilon^{+}\cup E^{-})$ then
12                   $\epsilon^{-}\leftarrow\texttt{getMCS}((\mathrm{KB}_{h}\cup\epsilon^{+})^{h}\cup(E^{-})^{s})$
13
14            return $\langle\epsilon^{+},\epsilon^{-}\rangle$
15
16       else
17             $\mathcal{C}\leftarrow\texttt{getMCS}(seed,\mathrm{KB}_{a}^{h}\cup\{\lnot\varphi^{h}\}\cup\mathrm{KB}_{\alpha}^{s})$
18             $\mathbb{R}\leftarrow\mathbb{R}\cup\{\mathcal{C}\}$
19
20
Algorithm 2 model-reconciling-explanation$(\mathrm{KB}_{\alpha},\mathrm{KB}_{a}^{h},\varphi)$
Table 2: Example of Algorithm 2 for computing a model reconciling explanation.
$\mathrm{KB}_{\alpha}=\;\{\stackrel{{\scriptstyle C_{1}}}{{(a\lor b)}},\stackrel{{\scriptstyle C_{2}}}{{(\lnot b\lor c)}},\stackrel{{\scriptstyle C_{3}}}{{\lnot c}},\stackrel{{\scriptstyle C_{4}}}{{(\lnot b\lor d)}},\stackrel{{\scriptstyle C_{5}}}{{\lnot d}}\}$ We have that $\mathrm{KB}_{\alpha}\models a$ and $\mathrm{KB}_{h}\not\models a$
$\mathrm{KB}_{h}=\;\{\stackrel{{\scriptstyle D_{1}}}{{b}},\stackrel{{\scriptstyle D_{2}}}{{\lnot c}}\}$
1. $\mathcal{R}\leftarrow\emptyset$
2. $\mathrm{KB}_{\alpha}^{h}\leftarrow\mathrm{KB}_{\alpha}\cap\mathrm{KB}_{h}=\{C_{3}\}$
3. $\mathrm{KB}_{\alpha}^{s}\leftarrow\mathrm{KB}_{\alpha}\setminus(\mathrm{KB}_{\alpha}\cap\mathrm{KB}_{h})=\{C_{1},C_{2},C_{4},C_{5}\}$
4. $E^{-}\leftarrow\{D_{1}\}$ MCS computed on $(\mathrm{KB}_{h}\setminus\mathrm{KB}_{\alpha})^{s}\cup\mathrm{KB}_{\alpha}^{h}$
5. $\mathrm{KB}_{h}\leftarrow\{D_{1},D_{2}\}\setminus\{D_{1}\}=\{D_{2}\}$
6. $seed\leftarrow\emptyset$ $minHS(\mathcal{R})$
7. $\{\lnot c\}\not\models a$ $SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
8. $\mathcal{C}\leftarrow\{C_{1}\}$ MCS computed on $\mathrm{KB}_{a}^{h}\cup\mathrm{KB}_{\alpha}^{s}\cup\{\lnot a^{h}\}$
9. $\mathcal{R}\leftarrow\{\{C_{1}\}\}$
10. $seed\leftarrow\{C_{1}\}$ $minHS(\mathcal{R})$
11. $\{\lnot c,a\vee b\}\not\models a$ $SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
12. $\mathcal{C}\leftarrow\{C_{2},C_{4}\}$ MCS computed on $\mathrm{KB}_{a}^{h}\cup\mathrm{KB}_{\alpha}^{s}\cup\{\lnot a^{h}\}$
13. $\mathcal{R}\leftarrow\{\{C_{1}\},\{C_{2},C_{4}\}\}$
14. $seed\leftarrow\{C_{1},C_{4}\}$ $minHS(\mathcal{R})$
15. $\{\lnot c,a\vee b,\lnot b\vee d\}\not\models a$ $SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
16. $\mathcal{C}\leftarrow\{C_{2},C_{5}\}$ MCS computed on $\mathrm{KB}_{a}^{h}\cup\mathrm{KB}_{\alpha}^{s}\cup\{\lnot a^{h}\}$
17. $\mathcal{R}\leftarrow\{\{C_{1}\},\{C_{2},C_{4}\},\{C_{2},C_{5}\}\}$
18. $seed\leftarrow\{C_{1},C_{2}\}$ $minHS(\mathcal{R})$
19. $\{\lnot c,a\vee b,\lnot b\vee c\}\models a$$\lnot SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
20. $\epsilon^{-}\leftarrow\{D_{1}\}$ MCS computed on $(\mathrm{KB}_{h}\cup\epsilon^{+})^{h}\cup(E^{-})^{s}$
21. $Return$$\langle\{C_{1},C_{2}\},\{D_{1}\}\rangle$ model reconciling explanation for $a$ from $\mathrm{KB}_{\alpha}$ for $\mathrm{KB}_{h}$
4.1.1 Model Reconciling Explanations
We have also previously showed how Algorithm 1 can be further extended for computing model reconciling explanations for an explanandum $\varphi$ from an agent knowledge base $\mathrm{KB}_{\alpha}$ for a human knowledge base $\mathrm{KB}_{h}$ (?). However, we only considered the specific task of finding a model reconciling explanation $\epsilon\subseteq\mathrm{KB}_{\alpha}\cup\mathrm{KB}_{h}$ such that $\mathrm{KB}_{h}\cup\epsilon\models\varphi$ and $\epsilon\setminus\mathrm{KB}_{h}$ is of minimum size. Notice that, in general, $\mathrm{KB}_{h}\cup\epsilon$ might be inconsistent. However, in our approach, we discard this possibility by preprocessing $\mathrm{KB}_{h}$. In particular, we create a new $\mathrm{KB}_{h}^{\prime}\subseteq\mathrm{KB}_{h}$ by removing a minimal set of formulae in $\mathrm{KB}_{h}$ that makes $\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha}$ inconsistent. The new $\mathrm{KB}_{h}^{\prime}$ is such that $\mathrm{KB}_{h}^{\prime}\cup\epsilon$ is always consistent.
We now modify this approach for computing model reconciling explanations $\mathcal{E}=\langle\epsilon^{+},\epsilon^{-}\rangle$, where $\epsilon^{+}\subseteq\mathrm{KB}_{\alpha}$ and $\epsilon^{-}\subseteq\mathrm{KB}_{h}$, such that $(\mathrm{KB}_{h}\cup\epsilon^{+})\setminus\epsilon^{-}\models\varphi$. Particularly, in addition to $\epsilon^{+}$, our approach now computes $\epsilon^{-}$ as well.
Algorithm 2 describes the pseudocode of our approach. At the beginning of the algorithm, we initialize $\mathbb{R}$ to the null set (line 2). $\mathbb{R}$ is used to store the MCSes, which acts as a mediator between $\mathrm{KB}_{\alpha}$ and $\mathrm{KB}_{h}$. Lines 2-2 are used to specify which clauses of $\mathrm{KB}_{\alpha}$ will be treated as hard and soft constraints, respectively. We then check if $\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha}$ is inconsistent (line 2). This is important in order to avoid the possibility of finding subsets $\epsilon^{+}$ that explain why $\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha}$ is inconsistent instead of the target explanandum. In case $\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha}$ is inconsistent, we preprocess $\mathrm{KB}_{h}$ by removing from $\mathrm{KB}_{h}\setminus\mathrm{KB}_{\alpha}$ a minimal set of formulae causing the conflict (i.e., an MCS) (lines 2-2), where $E^{-}$ stores the set of potential formulae $\epsilon^{-}$ to retract. The reconciliation procedure starts in line 2. The algorithm proceeds iteratively by computing a minimal hitting set on $\mathbb{R}$ and then testing for satisfiability the formulae $\epsilon^{+}$ (lines 2-2). The test checks whether adding $\epsilon^{+}$ to $\mathrm{KB}_{h}$ is sufficient for entailing $\varphi$. If $\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot\varphi\}$ is unsatisfiable, then $\mathrm{KB}_{h}\cup\epsilon^{+}\models\varphi$. In that case, the algorithm then checks whether $\mathrm{KB}_{h}\cup\epsilon^{+}\cup E^{-}$ is inconsistent, and if it is, it computes an MCS $e^{-}$ on $(\mathrm{KB}_{h}\cup\epsilon^{+})^{h}\cup(E^{-})^{s}$ (lines 2-2). The model reconciling explanation $\langle\epsilon^{+},\epsilon^{-}\rangle$ is then returned in line 2. Otherwise, the algorithm continues in line 2, where a new MCS is computed and added to $\mathbb{R}$ [^12] Note that the algorithm is complete as it is based on Algorithm 1, which is complete.
Table 2 shows an example trace of Algorithm 2.
4.2 Probabilistic Explanations
We now show how the algorithms described in the previous section can be used for computing probabilistic monolithic explanations (Definition 9) and probabilistic model reconciling explanations (Definition 14).
4.2.1 Monolithic Explanations
Consider an explanandum $\varphi$ and a belief base $\mathcal{B}$. First, notice that if we assume that the classical projection of $\mathcal{B}$ entails the explanandum $\varphi$, that is $\mathcal{B}^{\downarrow w}\models\varphi$, then Algorithm 1 can directly be applied on $\mathcal{B}^{\downarrow w}$ and $\varphi$ [^13] Recall that the classical projection of belief base $\mathcal{B}$ is the unweighted version of the set of formulae from $\mathcal{B}$. In that case, Algorithm 1 guarantees to find a monolithic explanation with maximum explanatory gain, since we know from Proposition 3 that explanatory gain achieves its maximum value for $\varphi$ when the monolithic explanation entails $\varphi$. Nevertheless, this does not guarantee that the monolithic explanation will be the most-preferred one, that is, the one with the highest explanatory power (Definition 13).
Obviously, a straightforward way of computing a most-preferred monolithic explanation is to use Algorithm 1 to enumerate all possible monolithic explanations for $\varphi$, and return the one that has the highest probability, which corresponds to the one with the highest explanatory power. But enumerating through all possible monolithic explanations and computing their probabilities can be computationally prohibited, as even extracting a smallest monolithic explanation is in $FP^{\Sigma^{p}_{2}}$ (?) and computing the probability of a formula is #$\#P$-complete (?, ?). We can, however, account for this high computational complexity by seeking for a monolithic explanation that is guaranteed to have a probability above a certain threshold.
First, the following lemma notes that for all possible monolithic explanations $\tilde{\epsilon}$ for explanandum $\varphi$, the following upper and lower probability bounds hold:
Lemma 2.
Let $\tilde{E}(\varphi)$ be the set of all monolithic explanations for explanandum $\varphi$ from belief base $\mathcal{B}$, where $\tilde{\epsilon}\models\varphi$ for all $\tilde{\epsilon}\in\tilde{E}(\varphi)$, and let $\omega_{1}$ be the most-probable world in which $\varphi$ is true. Then, for any $\tilde{\epsilon}\in\tilde{E}(\varphi)$, it holds that $P(\omega_{1})\leq P(\tilde{\epsilon})\leq P(\varphi)$.
Proof.
For the upper probability bound, since we assume that for all $\tilde{\epsilon}\in\tilde{E}(\varphi)$, $\tilde{\epsilon}\models\varphi$, then it must hold that for all $\tilde{\epsilon}\in\tilde{E}(\varphi)$, the worlds where $\tilde{\epsilon}$ is true are subsumed by the worlds where $\varphi$ is true (entailment property). This implies that for any $\tilde{\epsilon}\in\tilde{E}(\varphi)$, $P(\tilde{\epsilon})\leq P(\varphi)$.
For the lower bound, since $\omega_{1}$ is the most-probable world of $\varphi$, that is, the world where the highest number of formulae from $\mathcal{B}$ are satisfied, then all monolithic explanations for $\varphi$ must be true in $\omega_{1}$ (i.e., $\omega_{1}\models\tilde{\epsilon}$). As such, for any $\tilde{\epsilon}\in\tilde{E}(\varphi)$, $P(\tilde{\epsilon})\geq P(\omega_{1})$. ∎
However, some monolithic explanations may have a higher lower probability bound. Formally, we call such explanations $k$-bounded monolithic explanations:
Definition 19 ($k$-Bounded Monolithic Explanation).
[…]
Example 13.
Consider the belief base =$\mathcal{B}=\{(a,1),(\lnot a\vee b,3),(c,2),(\lnot c\vee b,1)\}$ and explanandum $b$. The two monolithic explanations for $b$ from $\mathcal{B}$ that entail $b$ are $\tilde{\epsilon}_{1}=\{a,\lnot a\vee b\}$ and $\tilde{\epsilon}_{2}=\{c,\lnot c\vee b\}$, where $P(\tilde{\epsilon}_{1})=0.64$ and $P(\tilde{\epsilon}_{2})=0.77$. Notice that there are four possible worlds in which $b$ is true: $\omega_{1}=\{a=T,b=T,c=T\},$ $\omega_{2}=\{a=F,b=T,c=T\}$, $\omega_{3}=\{a=T,b=T,c=F\}$, and $\omega_{4}=\{a=F,b=T,c=F\}$, where $P(\omega_{1})=0.57$, $P(\omega_{2})=0.20$, $P(\omega_{3})=0.07$, and $P(\omega_{4})=0.02$. The maximum number of intersections that entail $b$ is $k=2$ (i.e., $I_{2}=\{\lnot a\vee b,c,\lnot c\vee b\}$). Indeed, $\tilde{\epsilon}_{2}\subseteq I_{2}$ and $P(\tilde{\epsilon}_{2})=0.77=P(\omega_{1})+P(\omega_{2})$. Finally, notice how $\tilde{\epsilon}_{2}$ is also the most-preferred monolithic explanation for $b$ from $\mathcal{B}$; for $\gamma=0.5$, $\mathrm{EP}(\tilde{\epsilon}_{2},b)=0.57>\mathrm{EP}(\tilde{\epsilon}_{1},b)=0.50$.
Proposition 4.
Let $\mathcal{B}$ be a belief base and $\varphi$ an explanandum. A $1$-bounded monolithic explanation $\tilde{\epsilon}$ for $\varphi$ from $\mathcal{B}$ always exists.
Proof.
The proof follows directly from Lemma 2.
Interestingly, there also exists a maximal $k$-bounded monolithic explanation.
Corollary 2.
If $I_{k}\models\varphi$ and $I_{k+1}\not\models\varphi$, then $\exists\tilde{\epsilon}\subseteq I_{k}$ with maximal lower bound $P(\tilde{\epsilon})\geq P(I_{k})$
Proof.
[…]
The utility of a $k$-bounded monolithic explanation in computing probabilistic monolithic explanations can be described as follows. If we take the top $k$ most-probable worlds in which the explanandum $\varphi$ is true, then we can prune the search space of possible monolithic explanations by taking the intersection of those worlds and checking if it entails $\varphi$ – if it does, then we know that at least one monolithic explanation must be true in that world with probability at least equal to the sum of the probabilities of these top $k$ worlds. Building on this, we now present an algorithm for computing $k$-bounded monolithic explanations for $\varphi$ from $\mathcal{B}$, where we use Algorithm 1 as our core monolithic explanation generation engine.
1
Input: Belief base $\mathcal{B}$, explanandum $\varphi$, and user-defined parameter $\hat{k}$
Result: A $k$-bounded monolithic explanation $\tilde{\epsilon}$ for $\varphi$ from $\mathcal{B}$ for some $k\leq\hat{k}$
2
[…]
$\Omega_{\varphi}\leftarrow\texttt{getTopKWorlds}(\mathcal{B}\cup\{(\varphi,\infty)\},k)$
  // find candidate set of formulae
4 while true do
       // get intersecting formulae from top $k$ worlds of $\varphi$
5       $I_{k}\leftarrow\texttt{getIntersections}(\mathcal{B}^{\downarrow w},\Omega_{\varphi},k)$
6       if not SAT$(I_{k}\cup\{\lnot\varphi\})$ then
7             ϵ~~italic-$\tilde{\epsilon}\leftarrow$ monolithic-explanation$(I_{k},\varphi)$
8             return $\tilde{\epsilon}$
9
10       else
11             $k\leftarrow k-1$
12
Algorithm 3 probabilistic-monolithic-explanation$(\mathcal{B},\varphi,\hat{k})$
Algorithm 3 describes the main steps of our approach. The important factor is the user-defined parameter $\hat{k}$, which dictates the number of worlds of $\varphi$ to be considered. It is an integer with range $1\leq\hat{k}\leq|\Omega(\varphi)|$, where $\Omega(\varphi)$ is the set of all possible worlds of $\varphi$. Intuitively, the larger the $\hat{k}$, the more exhaustive the search will be as more worlds will be considered. The algorithm starts in line 1 with $k$ taking the user-defined value $\hat{k}$, and then proceeds to line 3, where it uses a weighted MaxSAT solver to find the top $k$ most-probable worlds of $\varphi$. Note that $(\varphi,\infty)$ denotes that $\varphi$ is added to the solver as a hard constraint. The main loop of the algorithm starts in line 3. In line 3, getIntersections extracts the set of intersecting formulae $I_{k}$ from $\mathcal{B}^{\downarrow w}$ that are true in worlds $\omega_{1}$ to $\omega_{k}$. If $I_{k}\models\varphi$, then we know that a monolithic explanation is in $I_{k}$ and the algorithm proceeds to use Algorithm 1 with $I_{k}$ and $\varphi$ as inputs to compute and return a monolithic explanation (lines 3-3). Otherwise, the algorithm discounts $k$ by $1$ and repeats the process until a suitable $k$ is found.
Algorithm 3 is complete in the sense that, eventually, a monolithic explanation will be returned.
Theorem 1.
Algorithm 3 is guaranteed to terminate with a solution.
Proof.
The proof rests on the fact that, in the worst case, the parameter $k$ will reach a value of $1$. This will then correspond to the most-probable world of $\varphi$, which entails all possible monolithic explanations for $\varphi$. From Lemma 2, we know that the most-probable world of $\varphi$ entails all possible monolithic explanations for $\varphi$, that is, for any $\tilde{E}(\varphi)$, $\omega_{1}\models\tilde{\epsilon}$, and $\tilde{\epsilon}\subseteq I_{1}$. Therefore, as Algorithm 3 uses $I_{1}$ as an input to Algorithm 1, which is guaranteed to return a solution, the algorithm is also guaranteed to terminate with a solution.
Theorem 2.
Algorithm 3 is guaranteed to return a maximal $k$-bounded monolithic explanation if the user-defined parameter $\hat{k}$ is initialized to $|\Omega(\varphi)|$.
Proof.
[…]
4.2.2 Model Reconciling Explanations
We now move on to the case of computing probabilistic model reconciling explanations ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for an explanandum $\varphi$ from an agent knowledge base $\mathrm{KB}_{\alpha}$ for a human belief base $\mathcal{B}_{h}$. Similarly to what we described in Section 4.2.1, Algorithm 2 can directly be used on $\mathrm{KB}_{\alpha}$ and $\mathcal{B}_{h}^{\downarrow w}$ for computing model reconciling explanations. Additionally, the concept of a $k$-bounded explanation (Definition 19) can also be used to guarantee a lower bound on the probability of $\tilde{\epsilon}^{+}$.
Algorithm 4 shows the pseudocode of our approach. The initial computational steps are similar to those in Algorithm 3, with the exception that $\mathrm{KB}_{\alpha}$ is now also considered in the computation of the most-probable worlds of the $\varphi$. Specifically, in line 4, $\mathrm{KB}_{\alpha}$ is converted into a belief base $\mathcal{B}_{\alpha}$ where each formula is given a weight that is larger than the sum of weights of $\mathcal{B}_{h}$. This is to enforce these formulae to be true in the worlds of the explanandum $\varphi$. Then, $\mathcal{B}_{\alpha}$ is used in conjunction with $\mathcal{B}_{h}$ to compute the top $k$ most-probable worlds of $\varphi$ (line 4). The algorithm proceeds in line 4 to extract formulae from $\mathrm{KB}_{\alpha}$ that are true in the first $k$ intersections of the worlds of $\varphi$. If they entail $\varphi$, the algorithm then proceeds to compute a model reconciling explanation by invoking Algorithm 2 (lines 4-4). Otherwise, the algorithm continues by discounting $k$ by $1$ and repeats the process.
Note that Algorithm 4 is complete and correct as it is based on Algorithms 2 and 3, which are complete and correct.
Input: Knowledge base $\mathrm{KB}_{\alpha}$, belief base $\mathcal{B}_{h}$, explanandum $\varphi$, and user-defined parameter $\hat{k}$
Result: A probabilistic model reconciling explanation ~$\tilde{\mathcal{E}}=\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$ for $\varphi$ from $\mathrm{KB}_{\alpha}$ for $\mathcal{B}_{h}$
1
[…]
3 $\mathrm{KB}_{\alpha}^{h}\leftarrow\mathrm{KB}_{\alpha}\cap\mathcal{B}_{h}^{\downarrow w}$
[…]
5 $\mathcal{B}_{\alpha}\leftarrow\{(\phi,W)\>|\>\phi\in\mathrm{KB}_{\alpha}\setminus\mathrm{KB}_{\alpha}^{h}\}$
6 $\Omega_{\varphi}\leftarrow\texttt{getTopKWorlds}(\mathcal{B}_{h}\cup\mathcal{B}_{\alpha}\cup\{(\varphi,\infty)\},k)$
7
8while $true$ do
9       $I_{k}\leftarrow\texttt{getIntersections}(\mathrm{KB}_{\alpha},\Omega_{\varphi},k)$
10       if not SAT$((I_{k}\cup\mathrm{KB}_{\alpha}^{h}\cup\{\lnot\varphi\}))$ then
11             $\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle\leftarrow\texttt{model-reconciling-explanation}(I_{k}\cup\mathrm{KB}_{\alpha}^{h},\mathcal{B}_{h}^{\downarrow w},\varphi)$
12             return $\langle\tilde{\epsilon}^{+},\tilde{\epsilon}^{-}\rangle$
13       else
14             $k\leftarrow k-1$
15
16
17
Algorithm 4 probabilistic-model-reconciling-explanation$(\mathrm{KB}_{\alpha},\mathcal{B}_{h},\varphi,\hat{k})$
5 Experimental Evaluations
This section presents a comprehensive evaluation of the proposed algorithms, assessing their effectiveness and efficiency across a range of scenarios.
5.1 Experimental Setup
Experiments were conducted on a system equipped with an M1 Max processor and 32GB of memory. The algorithms were implemented in Python, utilizing the PySAT toolkit (?) for SAT solving, MCS/MUS finding, weighted MaxSAT, and minimal hitting set computations. The time limit for all experiments was set to $500s$.
For our benchmarks, we selected a diverse set of problem instances:
  • $\bullet$
    Classical Planning Problems: We encoded classical planning problems from the International Planning Competition (IPC) in the style of ? (?), and used them as knowledge bases. The explanandum for each problem was the plan optimality query, which we constructed as described by ? (?).
  • $\bullet$
    Agent Scheduling Problems: We encoded logic-based agent scheduling problems based on the description by ? (?), and used them as the knowledge bases. The explanandum for each problem was a set of unsatisfied agent constraints.
  • $\bullet$
    Random CNF Problems: We generated random CNF formulae as knowledge bases using CNFgen (?). The explanandum for each problem was a conjunction of backbone literals, [^14] The backbone literals of a propositional knowledge base are the set of literals entailed by the knowledge base. which we computed using the minibones algorithm proposed by ? (?).
Note that we created associated belief bases for each problem by simply adding a random weight to each formula in the knowledge base. We will make the code and all benchmarks publicly available in the final version of the paper.
5.2 Results and Discussion
We now describe and discuss our experimental results, first for monolithic explanations and then for model reconciling explanations.
5.2.1 Monolithic Explanations
We evaluated Algorithm 3, referred to as alg3, on computing probabilistic monolithic explanations. Since the core monolithic explanation generation engine of alg3 is powered by Algorithm 1, referred to as alg1, we also evaluate its performance on the same instances. These experiments aim to answer the following questions:
  • Q1: What is the performance of alg3 on computing monolithic explanations across different problem instances?
  • Q2: Does the efficacy of alg3 change under different values of the user-defined parameter $\hat{k}$?
Parameter $\hat{k}$ Planning Scheduling Random CNF
S T/O Runtime S T/O Runtime S T/O Runtime
1 28 9 82.0s 30 5 80.0s 25 5 12.4s
50 32 5 79.0s 30 5 53.8s 25 5 8.6s
100 31 6 49.6s 30 5 44.7s 25 5 5.8s
150 31 6 45.5s 30 5 38.0s 25 5 3.4s
200 31 6 45.2s 30 5 37.2s 25 5 1.6s
Table 3: Number of Instances Solved (S) vs. Timed Out (T/O) by alg1 ($\hat{k}=1$) and alg3 ($\hat{k}=50$, $\hat{k}=100$, $\hat{k}=150$, $\hat{k}=200$).
(a) Planning Instances.
(b) Scheduling Instances.
(c) Random Instances.
Figure 1: Runtime distributions of alg1 ($\hat{k}=1$) and alg3 ($\hat{k}=50$, $\hat{k}=100$, $\hat{k}=150$, $\hat{k}=200$) across all planning, scheduling, and random CNF instances.
Table 3 tabulates the instances solved (i.e., found a monolithic explanation within the time limit) and not solved (i.e., timed out) by alg1 ($\hat{k}=1$) and alg3 at $\hat{k}=\{5,100,150,200\}$.151515alg3 at $\hat{k}=1$ corresponds to alg1 because each encoded knowledge base is consistent and entails the explanandum. As such, all formulae in the knowledge base are true in the most-probable world of the explanandum (i.e., $\hat{k}=1$), which means that alg3 reduces to alg1. We observe that the algorithm managed to solve most instances across different values of $\hat{k}$. Figure 1 shows the runtime distributions of alg3 across all values of $\hat{k}$ for computing a monolithic explanation. Interestingly, we observe that the runtimes decrease as $\hat{k}$ increases. This can be explained by the fact that for larger values of $\hat{k}$, alg3 considers the intersections of more worlds where the explanandum is true, which means that the number of formulae that are true in these intersections decreases. As such, the overall search space of monolithic explanations decreases as well, thus resulting in a reduced runtime needed for alg1 to compute a monolithic explanation. This can also be observed more granularly in Figure 2, where we can see the runtime distributions of alg1 ($\hat{k}=1$) and alg3 at $\hat{k}=200$ for each instance of the planning, scheduling, and random CNF problems. Again, the runtime of alg3 at $\hat{k}=200$ is smaller than that of alg1. Moreover, and as expected, in Figure 3, we can observe a positive correlation between runtime and the size of the encoded knowledge bases – as the size of the knowledge base increases, the runtimes increase as well. This is due to the fact that there is an increasing number of variables and formulae that must be considered, thus increasing the computational effort needed by the WMaxSAT, MCS, and hitting set solvers.
All of these observations indicate the feasibility and practical efficacy of alg3 across all benchmarks. In particular, from these experiments, we may conclude that the performance of alg3 increases as the user-defined parameter $\hat{k}$ increases. To reiterate, this is mainly because the overall search space of monolithic explanations that needs to be considered by alg1 (the main monolithic explanation generation engine) decreases. Finally, it is important to note that the performance of these algorithms lies in the effectiveness of the underlying WMaxSAT, MCS, and hitting set solvers. In other words, this also implies that any advancement in those solvers will automatically reflect in performance gains in our algorithms. Thus, future work can look at efficient and optimized solvers and examine whether there is any variability in performance.
(a) Planning Instances.
(b) Scheduling Instances.
(c) Random Instances.
Figure 2: Runtime distributions of alg1 ($\hat{k}=1$) and alg3 ($\hat{k}=200$) across commonly solved planning, scheduling, and random CNF instances.
(a) Planning Instances.
(b) Scheduling Instances.
(c) Random Instances.
Figure 3: Average runtime of alg1 ($\hat{k}=1)$ and alg3 ($\hat{k}=200$) to compute an explanation across different knowledge base sizes for the planning, scheduling, and random CNF instances.
5.2.2 Model Reconciling Explanations
We now examine the effectiveness of Algorithm 2, referred to as alg2, and Algorithm 4, referred to as alg4, on computing model reconciling explanations. We chose the value of $\hat{k}=200$ for alg4 as it was the better performing parameter for alg3 in our previous experiments. More specifically now, we are interested in scenarios with varying degrees of knowledge asymmetry between the agent and human models. To simulate such scenarios, we used the actual encoded knowledge bases as the model of the agent ($\mathrm{KB}_{\alpha}$), and tweaked that model and assigned it to be the model of the human ($\mathrm{KB}_{h}$ or $\mathcal{B}_{h}$). We considered the following ways to tweak the human model, resulting in the following five scenarios:
  • $\bullet$
    Scenario 1: We randomly removed 10% of the formulae and removed 20% of literals from 10% of the total formulae in the human’s model.
  • $\bullet$
    Scenario 2: We randomly removed 20% of the formulae and removed 20% of literals from 20% of the total formulae in the human’s model.
  • $\bullet$
    Scenario 3: We randomly removed 30% of the formulae and removed 20% of literals from 30% of the total formulae in the human’s model.
  • $\bullet$
    Scenario 4: We randomly removed 40% of the formulae and removed 20% of literals from 40% of the total formulae in the human’s model.
  • $\bullet$
    Scenario 5: We randomly removed 50% of the formulae and removed 20% of literals from 50% of the total formulae in the human’s model.
In general, these experiments aim to answer the following two questions:
  • Q1: What is the performance of the algorithms on computing model reconciling explanations across different problem instances?
  • Q2: What is the performance of the algorithms in scenarios with varying degrees of knowledge asymmetry between the agent and the human model?
Sce- Planning Scheduling Random CNF
nario alg2 alg4 alg2 alg4 alg2 alg4
S T/O Runtime S T/O Runtime S T/O Runtime S T/O Runtime S T/O Runtime S T/O Runtime
1 25 12 67.0s 28 9 59.7s 33 2 51.4s 33 2 33.1s 27 5 30.4s 21 11 12.3s
2 25 13 69.2s 27 10 71.8s 31 4 40.9s 31 4 28.0s 26 6 18.7s 20 12 0.5s
3 24 14 67.9s 26 12 68.8s 32 3 60.3s 32 3 37.1s 29 3 20.4s 21 11 2.5s
4 25 13 82.6s 27 11 84.0s 30 4 35.7s 30 4 22.9s 23 9 5.2s 20 12 0.5s
5 22 15 84.3s 24 13 89.9s 30 4 34.4s 30 4 21.5s 24 8 3.8s 20 11 0.5s
Table 4: Instances Solved (S) vs. Timed Out (T/O) for the Planning, Scheduling, and Random CNF Benchmarks for alg2 and alg4 at $\hat{k}=200$.
(a) Planning Instances.
(b) Scheduling Instances.
(c) Random Instances.
Figure 4: Runtime distributions of alg2 and alg4 at $\hat{k}=200$ to compute an explanation across all commonly solved instances.
(a) Planning Instances.
(b) Scheduling Instances.
(c) Random Instances.
Figure 5: Runtime distributions of alg2 and alg4 at $\hat{k}=200$ to compute an explanation across commonly solved instances in each of the five scenarios.
(a) alg2 on Planning Instances.
(b) alg4 on Planning Instances.
(c) alg2 on Scheduling Instances.
(d) alg4 on Scheduling Instances.
(e) alg2 on Random CNF Instances.
(f) alg4 on Random CNF Instances.
Figure 6: Distributions of the lengths of explanations $\epsilon^{+}$ and $\epsilon^{-}$ computed by alg2 and alg4 at $k=200$ across all planning, scheduling, and random CNF instances.
Table 4 tabulates the instances solved and timed out by alg2 and alg4 at $\hat{k}=200$ across the five scenarios, where we observe the following trends. For the planning instances, the runtime of both algorithms increases as the difference between the models of the agent and human increases (Scenarios 1 to 5), since both algorithms search over the explanation search space, which increases as the number of differences between the two models increases. As in the previous experiments, alg4 at $\hat{k}=200$ yields faster runtimes than alg2. For the scheduling instances, we observe that the runtimes increase from Scenario 1 to 3, but decrease from Scenarios 4 to 5. Upon closer inspection, this is mainly because the instances solved in these scenarios were easier (i.e., smaller knowledge base sizes) than those solved in the other three scenarios, thus resulting in smaller average runtimes. A similar trend is observed for the random CNF instances. However, in the random CNF instances, alg2 managed to solve more instances than alg4. After examining them more closely, we found that the main bottleneck of alg4 in those instances was computing the most-probable worlds of the explanandum (i.e., the WMaxSAT solver). Even for smaller values of $\hat{k}$, the solver failed to compute all the worlds under the specified time limit – the increase in search space (e.g., because of considering $\mathcal{B}_{\alpha}$ and $\mathcal{B}_{h}$) increased the complexity of these instances. We expect that an optimized and more dedicated solver may be able to overcome this limitation. The runtime distributions for alg2 and alg4 across all commonly solved instances and across commonly solved instances in each scenario can be seen in Figures 4 and 5, respectively. For these instances, we observe, like in the previous experiments, that alg4 has faster runtimes than alg2.
Moreover, in Figure 6 we see the distributions of the model reconciling explanation lengths computed by both algorithms. As expected, the general trend is that the size of the explanation $\epsilon^{+}$ (i.e., formulae from $\mathrm{KB}_{\alpha}$ for $\mathrm{KB}_{h}$ (or $\mathcal{B}_{h}$)) increases with each scenario, as the difference between the agent and human models increase. The same trend can be seen for $\epsilon^{-}$ – each scenario from 1 to 5 has an increasing amount of inconsistencies between the two models. Interestingly, $\epsilon^{-}$ was largest in the random CNF instances. This indicates that the inconsistencies between the human model and the corresponding $\epsilon^{+}$ were high. That can also be used to explain why alg4 failed to solve a subset of random CNF instances – highly inconsistent knowledge bases are considered as the most difficult instances for MaxSAT solvers.
In conclusion, the comparative analysis of alg2 and alg4 at $\hat{k}=200$ across varied problem instances shows some trends in performance and computational complexity. The observed increase in runtime with the increase of differences between agent and human models underscores the direct relationship between model disparity and the explanation search space size. Notably, alg4 consistently outperforms alg2 in terms of runtime across most scenarios, except in certain random CNF instances where the computation of most-probable worlds becomes a bottleneck due to the limitations of the WMaxSAT solver. This highlights a potential area for further optimization and development of more efficient solvers. Furthermore, the analysis of model reconciling explanation lengths reveals an expected increase in inconsistency measures as the model differences widen, particularly highlighted in random CNF instances.
6 Related Work
We now provide a discussion of related work from the planning and knowledge representation and reasoning (KR) literature. We focus on these two areas as our approach is motivated by the model reconciliation problem introduced by the planning community and it bears some similarity to other logic-based approaches in KR.
6.1 Related Planning Work
We have briefly discussed the body of work within explainable AI planning (XAIP) on the topic of model reconciliation in Section 1 as well as situated our work within that literature. We will now discuss other XAIP work, such as contrastive explanations, which take the form of “Why not A (instead of B)?”, where A is an alternative (or foil) suggested by the human to a decision B proposed by the agent. Contrastive explanations have found applications in linear temporal logic systems (?), general epistemic accounts (?), multi-agent optimization problems (?), and in oversubscription planning (?). There have also been used to provide a taxonomy of user questions that often arise during interactive plan exploration (?), as well as towards creating user interfaces for decision support systems (?, ?). For a more thorough discussion of the related work from the XAIP literature, please refer to the discussion by ? (?).
6.2 Related KR Work
In this paper, we presented a framework for probabilistic explanation generation in monolithic and model reconciliation scenarios. In the monolithic case, our definition of a probabilistic monolithic explanation (Definition 9) may appear similar to what was proposed by ? (?). Nonetheless, an important distinction here is that ? is dealing with epistemic states that do not contain the explanandum, while we are dealing with belief bases that do contain the explanandum. We also define a different notion of explanatory power as well as present algorithms for computing explanations. ? (?) have also considered the problem of defining what constitutes a (monolithic) explanation in probabilistic system, however they focus on epistemic states defined over causal structures.
The notion of (monolithic) explanation has also been explored by the probabilistic logic programming (PLP) community (?, ?), a formalism that extends logic programming languages (i.e., Prolog) with probabilities. In PLP, explanations have been associated with possible worlds. For instance, the MPE (most probable explanation) task consists in finding the world with the highest probability given some evidence (?). However, a world does not show the chain of inferences of a given explanandum and, moreover, it is not minimal by definition, since it usually includes a (possibly large) number of probabilistic facts whose truth value is irrelevant for the explanandum. Another alternative consists in using the proof of an explanandum as an explanation (?), where one can associate a proof with a (minimal) partial world $\omega^{\prime}$ such that for all worlds $\omega\supseteq\omega^{\prime}$, the explanandum is true in $\omega$. In this case, one can easily ensure minimality, but even if the partial world contains no irrelevant facts, it is still not easy to determine the chain of inferences behind a given explanandum. ? (?) have tackled explanation generation in PLP from the perspective of weighted model counting and knowledge compilation.
In the model reconciliation setting, we have extended our previous work on the logic-based model reconciliation problem (?) to a probabilistic case (Definition 14) for capturing scenarios where the human model is uncertain. ? (?) proposed a method for generating explanations in the case of uncertain human models, however, their approach is limited to planning problems, and importantly, it does not quantify the uncertainty levels of the generated explanations, that is, there is no notion of probabilistic explanation. In contrast, the application of probabilistic explanations in the context of model reconciliation that we consider in this work is, to our knowledge, novel.
Finally, the algorithms presented in this paper are an extension of our previous work (?). Specifically, Algorithms 1 and 2 are inspired by a procedure for computing a smallest minimal unsatisfiable set (SMUS) of an inconsistent formula, originally presented by ? (?). The method is also related to other similar approaches for enumerating MUSes and minimal correction sets (MCSes). Moreover, our approach is similar in spirit to the HS-tree presented by ? (?). Although the original purpose was to enumerate diagnoses, Reiter’s procedure can be easily adapted to enumerate MUSes (called conflicts in that paper) as already noted by ? (?). However, the computation of an SMUS might require more substantial modifications. Procedures like the one presented by Reiter, which target MCSes (diagnoses) instead of MUSes (conflicts), can be seen as the dual version of our algorithm. In particular, the algorithm MaxHS (?) applies the same idea of iteratively computing and testing a minimal hitting set for the computation of a MaxSAT solution (the complement of the smallest MCSes). Finally, there are other approaches that exploit the duality between MUSes and MCSes, but instead of iteratively checking if the current hitting set is an MUS, they first compute the set of all MCSes (?). This has the potential advantage that once all the MCSes are known, every minimal hitting set on the collection of all MCSes is guaranteed to be an MUS (Proposition 1). However, as the number of MCSes is, in the worst case, exponential in the size of the formula, this approach might fail even before reporting the first MUS. This is particularly unnecessary when the target is to return a single support, like the one presented in this paper.
7 Concluding Remarks
In this paper, we started by describing a framework for generating probabilistic monolithic explanations within uncertain knowledge bases (e.g., belief bases). Our work is an attempt to bridge the gap between classical explanation models and the inherent uncertainty that characterize real-world scenarios. We introduced the concepts of explanatory gain and explanatory power as quantitative measures to evaluate the effectiveness and relevance of explanations, thereby offering a better understanding of explanation quality.
Additionally, we presented an extension to the model reconciliation problem for generating probabilistic model reconciling explanations, which addresses the crucial need for reconciling model differences between an agent and a human model, specifically when the human model is not known with certainty. While we considered a classical knowledge base as the model of the agent in this work, future work could look into the assumption that the agent model is also probabilistic (i.e., a belief base). Reconciling two belief bases poses a significant challenge due to the complexity inherent in probabilistic logic and the need to account for uncertainty in both models. This task involves aligning two sets of beliefs, each associated with its own probability distributions, to achieve a coherent understanding that reflects the true state of affairs or intentions accurately. The process must consider not only the logical consistency between the belief bases but also the probabilistic coherence, ensuring that the reconciled belief base maintains a meaningful probability distribution over its assertions.
Furthermore, we developed algorithms that leverage the duality between minimal correction sets (MCSes) and minimal unsatisfiable subsets (MUSes) and demonstrated their potential for generating probabilistic explanations. Our experimental evaluations across diverse benchmarks underscore the practicality and effectiveness of these algorithms, validating their applicability in real-world settings. While our algorithms are presented within the framework of propositional logic, their underlying principles are broadly applicable to any constraint system where the satisfiability of subsets can be decided. This versatility opens avenues for extending our work to other logical systems, such as Markov Logic Networks (?) and Probabilistic Logic Programs (?).